Four numbers have a total of 92. The first number is twice the second number and the third number is 4 more than the sum of the first two. The fourth number is 3 less than the difference of the first two numbers. What are the numbers?
Answer by Earlsdon(6294)

Let the four numbers be a, b, c, and d.
From the problem description you can write:
1) a+b+c+d = 92 "The four numbers have a total of 92"
2) a = 2b "The first number is twice the second number..."
3) c = (a+b)+4 "...the third number is 4 more than the sum of the first two."
4) d = (a-b)-3 "The fourth number is 3 less than the difference of the first two numbers."
Substitute equation 2) into equation 1) to get:
1a) 2b+b+c+d = 92 Simplify.
2a) 3b+c+d = 92 Now substitute equation 3 into this equation.
2b) 3b+(a+b)+4+d = 92 Simplify this.
2b) a+4b+d+4 = 92 Substitute equation 2) into this equation.
2c) 2b+4b+d+4 = 94 Simplify.
2c) 6b+d+4 = 92 Substitute equation 4) into this equation.
2d) 6b+((a-b)-3)+4 = 92 Substitute equation 2) into this equation.
2e) 6b+((2b-b)-3)+4 = 92 Simplify.
2e) 7b+1 = 92 Subtract 1 from both sides.
7b = 91 Divide both sides by 7.
b = 13
a = 2b = 2(13) = 26
c = (a+b)+4 = 26+13+4 = 43
d = (a-b)-3 = 26-13-3 = 10
The numbers are:
a = 26
b = 13
c = 43
d = 10